One piece of copper jewelry at 117°C has exactly twice the mass of another piece, which is at 26.0°C. Both pieces are placed inside a calorimeter whose heat capacity is negligible. What is the final temperature inside the calorimeter (c of copper = 0.387 J/gK)?

thank you.
Answer #1

Let us assume mass of one copper jewelry is m g which is at 26 oC.

So, the mass of the copper jewlry at 117 oC is 2m g.

Let us assume the final temperature is x oC.

From the Law of Calorimetry the heat lost by the hot object = the heat gained by the cold object

Now, Heat lost = 2m * (117-x) * 0.387

        Heat gained = m * (x-26) * 0.387

So, we have 2m * (117-x) * 0.387 = m * (x-26) * 0.387

Solving we get, x = 86.7

So, the final temperature is 86.7 oC.